OpenStudy (anonymous):

Steven wants to start an IRA that will have $375,263 in it when he retires in 30 years. How much should he invest quarterly in his IRA to do this if the interest is 5.5% compounded semiannually? Assume an Annuity Due. Round to the nearest cent.

OpenStudy (anonymous):

An annuity due assumes the beginning balance to be P, ordinary annuity assumes the beginning balance is 0 ... from the other one we tried :) \[A=Pk^n+P\frac{1-k^n}{1-k}\] \[A=P(k^n+\frac{1-k^n}{1-k})\] \[\frac{A}{k^n+\frac{1-k^n}{1-k}}=P\] in this case, n = 30*4, and k = 1+.055/4